# Generic graded contractions of Lie algebras

### v1

@@ 5 Graded contractions as lax monoidal functors
Now we will consider the way to formalize the “generic” point of view on contractions using the functors introduced by Moody and Patera in [28], which we will define in a more general setting of lax monoidal functors on the category ${\rm Vec}^{G}$ of $G$-graded vector spaces. As noted in Section 2.2, there are different types of equivalences for graded contractions, two of which correspond to isomorphisms in ${\rm Vec}^{G}$, namely, strong equivalence (Definition 10) and equivalence via normalization (Definition 11). One of them implies the other, and Weimar-Woods conjectured that they, in fact, coincide:
@@
Two graded contractions $\varepsilon$ and $\varepsilon^{\prime}$ are strongly equivalent, $\varepsilon\approx\varepsilon^{\prime}$, if and only if they are equivalent via normalization, $\varepsilon\sim_{\rm n}\varepsilon^{\prime}$.
Draper et al. recently showed that this conjecture holds for the grading group $\mathbb{Z}_{2}^{3}$, by considering a special $\mathbb{Z}_{2}^{3}$-grading on the simple Lie algebra of type $G_{2}$, see discussion in [9, Section 1]. Note that, if we assume that the isomorphism in Definition 10 holds for a specific $G$-graded Lie algebra $\mathfrak{L}$, rather than for all Lie algebras, then this statement is false, even if $\mathfrak{L}$ is semisimple and the support of the grading is generates $G$, as discussed in [9, Example 2.12]. Since there exists a $G$-graded Lie algebra with the second support $\mathcal{P}_{G}$ (see Proposition 7), it is clear that $\varepsilon\approx\varepsilon^{\prime}$ implies $S(\varepsilon)=S(\varepsilon^{\prime})$.
@@
For the group $\mathbb{Z}$, if $\varepsilon\approx\varepsilon^{\prime}$, then there exists $\lambda\colon\mathbb{Z}\to\mathbb{F}^{\times}$ such that
$\varepsilon^{\prime}(m,n)=\varepsilon(m,n)(d\lambda)(m,n)\quad\mbox{for all }m\neq n.$
@@
$[L_{m},L_{n}]=(m-n)L_{m+n},\quad\forall m,n\in\mathbb{Z}.$
@@
Let $\varepsilon$ and $\varepsilon^{\prime}$ be strongly equivalent $\mathbb{Z}$-graded contractions of $\mathcal{W}$. This means that there exists an isomorphism of graded Lie algebras, $\varphi\colon\mathcal{W}^{\varepsilon}\to\mathcal{W}^{\varepsilon^{\prime}}$. Since $\dim\mathcal{W}_{n}=1$, for any $n\in\mathbb{Z}$ there exists $\lambda_{n}\in\mathbb{F}^{\times}$ such that $\varphi(L_{n})=\lambda_{n}L_{n}$. The equation $\varphi([L_{m},L_{n}]_{\varepsilon})=[\varphi(L_{m}),\varphi(L_{n})]_{\varepsilon^{\prime}}$ then gives the result. ∎
It is natural to consider Weimar–Woods conjecture within the categorical framework, where the notion of isomorphism is replaced with that of natural isomorphism. In fact, Moody and Patera initiated such consideration in [28, Section 4] even before Conjecture 31 appeared in the literature. They associated to graded contractions some specific endofunctors of the category of $G$-graded Lie algebras and further observed that the functors corresponding to contractions that are equivalent via normalization are naturally isomorphic.
@@
Lax monoidal functors. Here and in what follows we adopt the notation from [14]: a monoidal category is given by a quintuple $(\mathcal{C},\otimes,\mathsf{1},\alpha,\iota)$, where $\mathcal{C}$ is a category, $\otimes$ denotes the tensor product bifunctor, the natural isomorphism $\alpha$ determines the associativity constraint, $\mathsf{1}\in\mathcal{C}$ is the tensor unit, and $\iota\colon\mathsf{1}\otimes\mathsf{1}\to\mathsf{1}$ is an isomorphism, which satisfy certain axioms. When the context is clear, we may refer to this quintuple simply as $\mathcal{C}$. Furthermore, the isomorphism $\iota$ gives rise to the unit constraints $r$ and $l$ [14, Section 2.2], which may be used for an alternative difinition of a monoidal category [14, Definition 2.2.8].
@@
Let $(\mathcal{C},\otimes,\mathsf{1},\alpha,\iota)$ and $(\mathcal{C}^{\prime},\otimes^{\prime},\mathsf{1}^{\prime},\alpha^{\prime},\iota^{\prime})$ be monoidal categories. A lax monoidal functor between them is a triple $(F,J,\epsilon)$, where $F$ functor $F\colon\mathcal{C}\to\mathcal{C}^{\prime}$ together with coherence morphisms
@@
    a natural transformation $J(X,Y)\colon F(X)\otimes^{\prime}F(Y)\to F(X\otimes Y)$
@@
[…]
@@
A natural transformation between two lax monoidal functors is just a natural transformation repsecting the monoidal structure.
@@
Let $(F,J,\epsilon)$ and $(G,I,\kappa)$ be lax monoidal functors from the monoidal category $(\mathcal{C},\otimes,\mathsf{1},\alpha,\iota)$ to the monoidal category $(\mathcal{C}^{\prime},\otimes^{\prime},\mathsf{1}^{\prime},\alpha^{\prime},\iota^{\prime})$. A monoidal natural transformation $\nu$ from $(F,J,\epsilon)$ to $(G,I,\kappa)$ is a natural transformation $\nu_{X}\colon F(X)\to G(X)$ between the underlying functors that is compatible with the tensor product and the unit in that the following diagrams commute for all objects $X,Y\in\mathcal{C}$:
[…]
@@
The identity endofunctor $\mathrm{Id}\colon{\rm Vec}^{G}\to{\rm Vec}^{G}$ can be given the structure of a monoidal functor as follows. Given a function $\sigma\colon G\times G\to\mathbb{F}^{\times}$, we take as $\varepsilon$ the identity of $\mathsf{1}$ define a natural isomorphism $J^{\sigma}_{V,W}\colon V\otimes W\to V\otimes W$,
@@
$\beta^{\prime}(g,h)\sigma(h,g)=\beta(g,h)\sigma(g,h)\quad\mbox{for all}\ g,h\in G.$
@@
The following result shows that Conjecture 31 holds when the strong equivalence of contractions is replaced with “natural” strong equivalence.
@@
The lax monoidal functors $F_{\sigma}=(\mathrm{Id},J_{\sigma})$ and $F_{\sigma^{\prime}}=(\mathrm{Id},J_{\sigma^{\prime}})$ are naturally isomorphic if and only if $\varepsilon\sim_{\rm n}\varepsilon^{\prime}$, i.e., $\varepsilon^{\prime}=\varepsilon\,d\alpha$ for some $\alpha\colon G\to\mathbb{F}^{\times}$.
@@
$\sigma(g,h)=\alpha(gh)^{-1}\alpha(g)\alpha(h)\sigma^{\prime}(g,h)$ (13)
@@
${V}$${W}$${V}$${W}$$\scriptstyle{f}$$\scriptstyle{\nu_{V}}$$\scriptstyle{\nu_{W}}$$\scriptstyle{f}$
commutes, hence $\nu$ is a natural isomorphism from the identity endofunctor to itself. We need to verify that $\nu$ respects the lax structure. The commutativity of the first diagram from Definition 34, applied to $F_{\sigma}$ and $F_{\sigma^{\prime}}$,
@@
Conversely, suppose we have a natural isomorphism $\nu\colon F_{\sigma}\to F_{\sigma^{\prime}}$ of lax monoidal functors. Consider the simple objects $\delta_{g}$ in ${\rm Vec}^{G}$. Since $\nu_{\delta_{g}}\colon\delta_{g}\to\delta_{g}$ is an isomorphism, there exists $\alpha(g)\in\mathbb{F}^{\times}$ such that $\nu_{\delta_{g}}(v)=\alpha(g)v$ for all $v\in\delta_{g}$. In view of this, the commutativity of the diagram
[…]
@@ 6 Conclusion
More specifically, Moody and Patera studied graded contractions from the perspectives of semigroup cohomology and endofunctors of the category of Lie algebras, see more details in Sections 3 and 5. Weimar-Woods, on the other hand, developed a theory based on the properties of what she called “surviving defining relations” and “surviving higher-order identities”, seemingly unrelated to cohomology. This led to a classification theorem for generic graded contractions [39, Theorem 7.1] over the fields of complex and real numbers.
@@
Now we summarize the results of this paper. We began Section 3 with establishing in Corollary 17 that the group $\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$ is a free abelian group of rank $\leqslant|G|$ using an analogue of the bar resolution. In Remark 18, we discussed the freeness of the group $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ claimed in [39, Theorem 6.5], whose proof appears to be incorrect. Nevertheless, the numerical experiments carried out for low-order grading groups lead us to pose the following question:
@@
Then we introduced the abelian group $H^{2}_{S}(A)={\rm Hom}(\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle,A)/B^{2}_{S}(A)$, which in the case $A=\mathbb{F}^{\times}$ and $S\in\mathcal{S}(G)$ classifies the $\mathbb{F}$-valued contractions with support $S$ up to equivalence via normalization. In Theorem 19, we presented exact sequences that allow one to find the group $H^{2}_{S}(A)$. The special cases when $\mathbb{F}$ is an algebraically closed or a real closed field were presented in Corollaries 21 and 22, respectively. These corollaries were related to Theorem 7.1 from [39] in Remark 23, thereby providing a cohomological interpretation of the notion of sign invariants from [39].
Although we established the classification theorem (Theorem 19) for generic graded contractions, there is still one important question to address:
For any finite group $G$, develop an efficient algorithm for finding the set of supports of generic graded contractions, $\mathcal{S}(G)$.
We touched upon this problem in Remark 13 and in Section 4.
In Section 4, we studied the affine algebraic variety of $G$-graded contractions and its cellular decomposition. Using this decomposition, in Theorem 27 we described the vanishing ideal of the closure of the ${\rm T}_{S}$-orbit of an arbitrary generic graded contraction with support $S$. This allowed us to relate graded contractions to graded degenerations, introduced in Section 2.1, and to Inönü–Wigner contractions.
We revisited Conjecture 31 from [39] in Section 5. Interpreting generic $G$-graded contractions as lax monoidal structures on the identity endofunctor of the monoidal category of $G$-graded vector spaces, we establish a functorial version of this conjecture.
Acknowledgements. SDK is grateful to Dr. Alex Bihlo for the support of his research thanks to funding from the Canada Research Chairs program, the InnovateNL Leverage R&D Program and the NSERC Discovery Grant. MVK is supported by NSERC Discovery Grant 2018-04883.

### v2

@@ 5 Graded contractions as lax monoidal functors
Now we will consider the way to formalize the “generic” point of view on contractions using the functors introduced by Moody and Patera in [36], which we will define in a more general setting of lax monoidal functors on the category ${\rm Vec}^{G}$ of $G$-graded vector spaces. As noted in Section 2.3, there are different types of equivalences for graded contractions, two of which correspond to isomorphism in ${\rm Vec}^{G}$, namely, isomorphism (Definition 10) and equivalence via normalization (Definition 11). One of them implies the other, and Weimar–Woods conjectured that they, in fact, coincide:
@@
Two graded contractions $\varepsilon$ and $\varepsilon^{\prime}$ are isomorphic, $\varepsilon\simeq\varepsilon^{\prime}$, if and only if they are equivalent via normalization, $\varepsilon\sim_{\rm n}\varepsilon^{\prime}$.
Draper et al. recently showed that this conjecture holds for the grading group $\mathbb{Z}_{2}^{3}$, by considering a special $\mathbb{Z}_{2}^{3}$-grading on the simple Lie algebra of type $G_{2}$, see discussion in [12, Section 1]. Note that, if we assume that the isomorphism in Definition 10 holds for a specific $G$-graded Lie algebra $\mathfrak{L}$, rather than for all Lie algebras, then this statement is false, even if $\mathfrak{L}$ is semisimple and the support of the grading generates $G$, as discussed in [12, Example 2.12]. Since there exists a $G$-graded Lie algebra with the second support $\mathcal{P}_{G}$ (see Proposition 7), it is clear that $\varepsilon\simeq\varepsilon^{\prime}$ implies $S(\varepsilon)=S(\varepsilon^{\prime})$.
@@
For the group $\mathbb{Z}$, if $\varepsilon\simeq\varepsilon^{\prime}$, then there exists $\lambda\colon\mathbb{Z}\to\mathbb{F}^{\times}$ such that
$\varepsilon^{\prime}(m,n)=\varepsilon(m,n)\,d\lambda(m,n)\quad\mbox{for all }m\neq n.$
@@
$L_{m},L_{n}]=(m-n)L_{m+n},\quad\forall m,n\in\mathbb{Z}.$
@@
Let $\varepsilon$ and $\varepsilon^{\prime}$ be isomorphic $\mathbb{Z}$-graded contractions of $\mathcal{W}$. This means that there exists an isomorphism of graded Lie algebras, $\varphi\colon\mathcal{W}^{\varepsilon}\to\mathcal{W}^{\varepsilon^{\prime}}$. Since $\dim\mathcal{W}_{n}=1$, for any $n\in\mathbb{Z}$ there exists $\lambda_{n}\in\mathbb{F}^{\times}$ such that $\varphi(L_{n})=\lambda_{n}L_{n}$. The equation $\varphi([L_{m},L_{n}]_{\varepsilon})=[\varphi(L_{m}),\varphi(L_{n})]_{\varepsilon^{\prime}}$ then gives the result. ∎
It is natural to consider Weimar–Woods conjecture within the categorical framework, where the notion of isomorphism is replaced with that of natural isomorphism. In fact, Moody and Patera initiated such consideration in [36, Section 4] even before Conjecture 31 appeared in the literature. They associated to graded contractions some specific endofunctors of the category of $G$-graded Lie algebras and further observed that the functors corresponding to contractions that are equivalent via normalization are naturally isomorphic.
@@
Lax monoidal functors. Here and in what follows we adopt the notation from [18]: a monoidal category is given by a quintuple $(\mathcal{C},\otimes,\mathsf{1},\alpha,\iota)$, where $\mathcal{C}$ is a category, $\otimes$ denotes the tensor product bifunctor, the natural isomorphism $\alpha$ determines the associativity constraint, $\mathsf{1}\in\mathcal{C}$ is the tensor unit, and $\iota\colon\mathsf{1}\otimes\mathsf{1}\to\mathsf{1}$ is an isomorphism, which satisfy certain axioms. When the context is clear, we may refer to this quintuple simply as $\mathcal{C}$. Furthermore, the isomorphism $\iota$ gives rise to the unit constraints $r$ and $l$ [18, Section 2.2], which may be used for an alternative definition of a monoidal category [18, Definition 2.2.8].
@@
Let $(\mathcal{C},\otimes,\mathsf{1},\alpha,\iota)$ and $(\mathcal{C}^{\prime},\otimes^{\prime},\mathsf{1}^{\prime},\alpha^{\prime},\iota^{\prime})$ be monoidal categories. A lax monoidal functor between them is a triple $(F,J,\epsilon)$, where $F$ is a functor $F\colon\mathcal{C}\to\mathcal{C}^{\prime}$ together with coherence morphisms
@@
    a natural transformation $J_{X,Y}\colon F(X)\otimes^{\prime}F(Y)\to F(X\otimes Y)$
@@
    $\big(F(X)\otimes^{\prime}F(Y)\big)\otimes^{\prime}F(Z)$$F(X)\otimes^{\prime}\big(F(Y)\otimes^{\prime}F(Z)\big)$$F(X\otimes Y)\otimes^{\prime}F(Z)$$F(X)\otimes^{\prime}F(Y\otimes Z)$$F\big((X\otimes Y)\otimes Z\big)$$F\big(X\otimes(Y\otimes Z)\big)$$\scriptstyle{\alpha^{\prime}_{F(X),F(Y),F(Z)}}\scriptstyle{\simeq}$$\scriptstyle{{J_{X,Y}}\otimes^{\prime}{\rm id}}$$\scriptstyle{{\rm id}\otimes^{\prime}{J_{Y,Z}}}$$\scriptstyle{J_{X\otimes Y,Z}}$$\scriptstyle{J_{X,Y\otimes Z}}$$\scriptstyle{F(\alpha_{X,Y,Z})}\scriptstyle{\simeq}$ (12)
@@
A natural transformation between two lax monoidal functors is just a natural transformation respecting the monoidal structure.
@@
Let $(F,J,\epsilon)$ and $(G,I,\kappa)$ be lax monoidal functors from the monoidal category $(\mathcal{C},\otimes,\mathsf{1},\alpha,\iota)$ to the monoidal category $(\mathcal{C}^{\prime},\otimes^{\prime},\mathsf{1}^{\prime},\alpha^{\prime},\iota^{\prime})$. A monoidal natural transformation $\nu$ from $(F,J,\epsilon)$ to $(G,I,\kappa)$ is a natural transformation $\nu_{X}\colon F(X)\to G(X)$ between the underlying functors that is compatible with the tensor products and the units in that categories, that is, the following diagrams commute for all objects $X,Y\in\mathcal{C}$:
$F(X)\otimes^{\prime}F(Y)$$G(x)\otimes^{\prime}G(Y)$$F(X\otimes Y)$$G(X\otimes Y)$$\scriptstyle{\nu_{X}\otimes^{\prime}\nu_{Y}}$$\scriptstyle{J_{X,Y}}$$\scriptstyle{I_{X,Y}}$$\scriptstyle{\nu_{X\otimes Y}}$  $\mathsf{1}^{\prime}$$F(\mathsf{1})$$G(\mathsf{1})$$\scriptstyle{\epsilon}$$\scriptstyle{\kappa}$$\scriptstyle{\nu_{1}}$
@@
The identity endofunctor $\mathrm{Id}\colon{\rm Vec}^{G}\to{\rm Vec}^{G}$ can be given the structure of a monoidal functor as follows. Given a function $\sigma\colon G\times G\to\mathbb{F}^{\times}$, we take as $\epsilon$ the identity of $\mathsf{1}$ define a natural isomorphism $J^{\sigma}_{V,W}\colon V\otimes W\to V\otimes W$,
@@
$\beta^{\prime}(g,h)\varepsilon(h,g)=\beta(g,h)\varepsilon(g,h)\quad\mbox{for all}\ g,h\in G.$
@@
The following result shows that Conjecture 31 holds when the isomorphism of contractions is replaced with natural isomorphism.
@@
The lax monoidal functors $F_{\varepsilon}:=(\mathrm{Id},J^{\varepsilon},{\rm id}_{\mathbb{F}})$ and $F_{\varepsilon^{\prime}}:=(\mathrm{Id},J^{\varepsilon^{\prime}},{\rm id}_{\mathbb{F}})$ are naturally isomorphic if and only if $\varepsilon\sim_{\rm n}\varepsilon^{\prime}$, i.e., $\varepsilon^{\prime}=\varepsilon\,d\alpha$ for some $\alpha\colon G\to\mathbb{F}^{\times}$.
@@
$\varepsilon(g,h)=\alpha(gh)^{-1}\alpha(g)\alpha(h)\varepsilon^{\prime}(g,h).$ (13)
@@
commutes, hence $\nu$ is a natural isomorphism from the identity endofunctor to itself. We need to verify that $\nu$ respects the lax structure. The commutativity of the first diagram from Definition 34, applied to $F_{\varepsilon}$ and $F_{\varepsilon^{\prime}}$,
$V\otimes W$$V\otimes W$$V\otimes W$$V\otimes W$$\scriptstyle{\nu_{V}\otimes\nu_{W}}$$\scriptstyle{J^{\varepsilon}_{V,W}}$$\scriptstyle{J^{\varepsilon^{\prime}}_{V,W}}$$\scriptstyle{\nu_{V\otimes W}}$
@@
Conversely, suppose we have a natural isomorphism $\nu\colon F_{\varepsilon}\to F_{\varepsilon^{\prime}}$ of lax monoidal functors. Consider the simple objects $\delta_{g}$ in ${\rm Vec}^{G}$. Since $\nu_{\delta_{g}}\colon\delta_{g}\to\delta_{g}$ is an isomorphism, there exists $\alpha(g)\in\mathbb{F}^{\times}$ such that $\nu_{\delta_{g}}(v)=\alpha(g)v$ for all $v\in\delta_{g}$. In view of this, the commutativity of the diagram
$\delta_{g}\otimes\delta_{h}$$\delta_{g}\otimes\delta_{h}$$\delta_{g}\otimes\delta_{h}$$\delta_{g}\otimes\delta_{h}$$\scriptstyle{\nu_{\delta_{g}}\otimes\nu_{\delta_{h}}}$$\scriptstyle{J^{\varepsilon}_{\delta_{g},\delta_{h}}}$$\scriptstyle{J^{\varepsilon^{\prime}}_{\delta_{g},\delta_{h}}}$$\scriptstyle{\nu_{\delta_{g}\otimes\delta_{h}}}$
@@ 6 Conclusion
More specifically, Moody and Patera studied graded contractions from the perspectives of semigroup cohomology and endofunctors of the category of Lie algebras, see more details in Sections 3 and 5. Weimar–Woods, on the other hand, developed a theory based on the properties of what she called “surviving defining relations” and “surviving higher-order identities”, seemingly unrelated to cohomology. This led to a classification theorem for generic graded contractions [48, Theorem 7.1] over the fields of complex and real numbers.
@@
Now we summarize the results of this paper and state some open problems. We introduced the abelian group ${\rm H}^{2}_{S}(A)={\rm Hom}(\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle,A)/{\rm B}^{2}_{S}(A)$, which in the case $A=\mathbb{F}^{\times}$ and $S\in\mathcal{S}(G)$ classifies the $\mathbb{F}$-valued contractions with support $S$ up to equivalence via normalization. In Theorem 19, we presented exact sequences that allow one to find the group ${\rm H}^{2}_{S}(A)$. The special cases when $\mathbb{F}$ is an algebraically closed or a real closed field were presented in Corollaries 21 and 22, respectively. These corollaries were related to Theorem 7.1 from [48] in Remark 23, thereby providing a cohomological interpretation of the notion of sign invariants from [48].
Although we established the classification theorem (Theorem 19) for generic graded contractions, there is still an important question to address:
Question 1.
Develop an efficient algorithm for finding the set of supports of generic graded contractions, $\mathcal{S}(G)$, for any finite abelian group $G$.
We touched upon this problem in Remark 13 and in Section 4.
We established in Corollary 17 that $\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$ is a free abelian group of rank $\leqslant|G|$. More generally, the freeness of the group $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ is claimed in [48, Theorem 6.5], but the proof appears to be incorrect (see Remark 18). Nevertheless, the numerical experiments that we carried out for low-order grading groups lead us to pose the following question:
Question 2.
@@
As mentioned above, there are many works studying graded contractions of a particular $G$-graded Lie algebra. We suggest studying generic graded contractions for particular grading groups or their classes instead of fixing the Lie algebra:
Question 3.
Describe the group ${\rm H}^{2}_{S}(\mathbb{F}^{\times})$ for certain groups or classes of abelian groups, e.g., cyclic groups, elementary $p$-groups, etc.
In Section 4, we studied the affine algebraic variety $X$ of $G$-graded contractions and its cellular decomposition. In Theorem 27, we described the vanishing ideal of the closure of the ${\rm T}_{S}$-orbit of an arbitrary generic graded contraction with support $S$. This allowed us to relate graded contractions to graded degenerations, introduced in Section 2.2, and to Inönü–Wigner contractions. We also described the irreducible components of the varieties of generic graded contractions of the grading groups $\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$, see Examples 25 and 26. This leads us to the following question:
Question 4.
How does the structure of the variety $X$ of generic graded contractions depend on the structure of the grading group $G$?
Since the ideal defining $X$ is generated by binomials, it follows from [16] that the vanishing ideals of $X$ and its irreducible components are also binomial, which implies that the irreducible components of $X$ are affine toric varieties, see, e.g., [9]. (A particular case is the irreducible component $X(\bar{1})$ consisting of continuous graded contractions, see Proposition 24.) Moreover, $X$ is defined by homogeneous polynomials, so we can consider the corresponding projective variety and its components. This observation paves the way to using a wide range of methods from toric geometry, such as combinatorics of polyhedra, for further study of the components of $X$.
Finally, in Section 5, we revisited Conjecture 31 from [48]. Interpreting generic $G$-graded contractions as lax monoidal structures on the identity endofunctor of the monoidal category of $G$-graded vector spaces, we established a functorial version of this conjecture, but the original remains open.
Acknowledgments
The authors are grateful to the anonymous reviewer whose feedback helped improve the exposition. MVK is supported by NSERC Discovery Grant 2018-04883. SDK is grateful to Dr. Alex Bihlo for the support of his research thanks to funding from the Canada Research Chairs program, the InnovateNL Leverage R&D Program and the NSERC Discovery Grant. SDK also acknowledges support from Dr. Yuri Bahturin’s NSERC Discovery Grant.